Apply gauss's theorem to calcutta the electric field in the case of a changed infimite plan
1
Expert's answer
2021-07-15T12:48:46-0400

Gaussian surface total output flux


"\phi=\smallint_{s_1}E.ds+\smallint_{s_2}E.ds+\smallint_{s_3}E.ds"

s3 surface Eand ds is perpendicular

"\theta=90°"

"\smallint_{s_3}E.ds=\smallint_{s_3}E.dscos\theta\\\smallint_{s_3}E.dscos90=0"

"s_1,s_3" Surface are E and ds are same direction

"\theta=0°"


"\phi=\smallint_{s_1}E.dscos\theta+0+\smallint_{s_3}E.dscos\theta\\\phi=2ES\rightarrow(1)"

Gauss law

"\phi=\frac{q}{\epsilon_0}\rightarrow(2)"

Surface charge density

"\sigma=\frac{q}{S}\\q=\sigma S \rightarrow(3)"

equation (1),(2),(3) we can written as

"E=\frac{\sigma}{2\epsilon}"